Magnetic moment of a bar magnet is M. the work done to turn the magnet by 90 degrees of magnet in direction of magnetic field B will be
 

Grade:12th pass

3 Answers

Eshan
5 years ago
Dear student,

Torque acting on a magnet of magentic moment\vec{M}kept in an external magnetic field\vec{B}is
\vec{\tau}=\vec{M}\times \vec{B}=MBsin\theta

Thus the work done in moving a magnet by90^{\circ}in direction of magnetic field is
W=\int_{90^{\circ}}^{0^{\circ}}MBsin\theta d\theta=MB(cos90^{\circ}-cos0^{\circ}) =-MB
Gitanjali Rout
5 years ago
W=\int_{90^{\circ}}^{0^{\circ}}MBsin\theta d\theta=MB(cos90^{\circ}-cos0^{\circ}) =-MBin direction of magnetic field is90^{\circ}Thus the work done in moving a magnet by\vec{\tau}=\vec{M}\times \vec{B}=MBsin\thetais\vec{B}kept in an external magnetic field\vec{M}orque acting on a magnet of magentic moment
Khimraj
5 years ago
Work done is given by
W = MB(cos0 – cos90) = MB.
Hope it clears.